OpenStudy (anonymous):

A store received $823 from the sale of 5 tape recorders and 7 radios. If the receipts from the tape recorders exceeded the receipts from the radios by $137, what was the cost of a tape recorder? i got 49..is that right??? (:

OpenStudy (amistre64):

5t + 7r = 823 t -137 = r 5t + 7(t-137) = 823 5t + 7t-(7)137) = 823 13t = 823 + 7(137) t = (823 + 7(137))/13 = 137.08/5 = 27.42 maybe? i might have a wrong interpretation in my equations tho

OpenStudy (amistre64):

5t = 7r + 137 ?? maybe

OpenStudy (amistre64):

14r = 823 - 137 r = 49 then.... 5t = 7(49) + 137 t = [7(49) + 137]/14 = 96 then

OpenStudy (amistre64):

t = [7(49) + 137]/5 = 96 ..... i have no idea where the 14 came from lol

OpenStudy (anonymous):

thanks(: